# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 2.3.4. Two-Stream Convolutional Network for Fake Content Detection by Yousaf: 2022
Yousaf et al. (Yousaf et al., 2022) proposed TwoStreamNet, a two-stream convolutional neural network designed to enhance the generalizability of fake visual content detection by jointly analyzing spatial and frequency features. The network comprises two main modules: the Spatial Stream and the Frequency Stream, which independently process input images and fuse their outputs at the classification stage to improve detection accuracy.
@@
The Frequency Stream captures frequency domain artifacts by first converting images to the YCbCr color space to decorrelate color channels. Discrete Fourier Transform (DFT) and Discrete Wavelet Transform (DWT) (see equation (7)) are applied to each channel. DFT decomposes image signals into real and imaginary components, while DWT captures both low- and high-frequency sub-bands. The resulting frequency features, represented as $H\times W\times 18$ feature maps, are processed using ResNet-50 (He et al., 2016) to extract discriminative frequency patterns. The Spatial Stream focuses on spatial domain features by processing original RGB images augmented with Gaussian blur and JPEG compression, similar to the approach in (Wang et al., 2020b). These augmented images are passed through a ResNet-50 network to extract spatial features. Finally, the outputs of the two streams are fused via probability averaging, ensuring equal contributions from spatial and frequency domains to the final decision. This combined framework highlights the importance of integrating frequency features for robust detection of synthesized visual content.
2.3.5. Synthbuster by Bammey: 2023
@@
Bammey (Bammey, 2024) introduced Synthbuster, a forensic method for detecting synthetic images generated by state-of-the-art diffusion models. The method begins by applying a cross-difference filter, originally defined by Chen et al. (Chen and Hsu, 2008), to highlight periodic frequency artifacts in synthetic images. The filter acts as a high-pass filter, generating residual images using the operation:
$C(x,y)=|I(x,y)+I(x+1,y+1)-I(x+1,y)-I(x,y+1)|,$
@@
The Fast Fourier Transform (FFT) is then applied to the residual image, extracting spectral components corresponding to periodicities $0,2,4,$ and $8$ in both vertical and horizontal directions. The resulting features form a 135-dimensional magnitude peak vector for the RGB channels. These features are classified using a histogram-based gradient boosting tree classifier (HBGB) (Ke et al., 2017), trained specifically to distinguish synthetic images from real ones.
@@
2.3.6. Meng’s Artifact Feature Purification: 2024
@@
Meng et al. (Meng et al., 2024) addressed two key limitations in existing detection methods: poor generalization across generative models and limited effectiveness on images from diverse large-scale datasets. To overcome these challenges, they proposed the Artifact Purification Network (APN), which extracts generalizable artifact features through explicit and implicit purification processes. Explicit purification isolates artifact features in spatial and frequency domains by employing feature decomposition and frequency-band proposals to detect suspicious patterns. Implicit purification, guided by a classifier, further refines these features using mutual information estimation, enhancing the robustness of detection across various generators and datasets. This dual purification approach significantly improves generalization and detection accuracy.
2.3.7. An Image Transformation Perspective by Li: 2024
@@
Li et al. (Li et al., 2024b) identified biases in existing detection methods, including weakened and overfitted artifact features, which limit their generalization capability. To address these challenges, the authors proposed the SAFE (Simple Preserved and Augmented Features) framework, designed to improve generalizable artifact detection by integrating effective transformations with preprocessing and augmentations. The framework employs several key strategies: (1) replacing conventional down-sampling with a crop operator (RandomCrop for training and CenterCrop for inference) to preserve local correlations and prevent artifact distortion; (2) applying invariant augmentations, such as ColorJitter and RandomRotation, to mitigate color discrepancies and irrelevant rotation-related features; (3) using a patch-based random masking strategy to enhance sensitivity to local image regions and subtle artifacts; and (4) incorporating Discrete Wavelet Transform (DWT) to extract high-frequency features critical for distinguishing synthetic from real images. The processed data was used to train a ResNet model (He et al., 2016) with 1.44M parameters, achieving strong detection accuracy during testing. This approach demonstrated the effectiveness of integrating diverse transformations to improve detection robustness and generalization.
2.3.8. Yan’s AIDE Method for Synthetic Image Detection (2024)
@@
Yan et al. (Yan et al., 2024) introduced the Chameleon dataset, a collection of AI-generated images designed to closely resemble real-world scenes, highlighting the limitations of existing detectors, which frequently misclassify these images as real. To address these challenges, the authors proposed AIDE (AI-generated Image Detector with Hybrid Features), a method that combines low-level frequency-based features with high-level semantic embeddings from CLIP. AIDE processes input images by dividing them into patches and applying discrete cosine transform (DCT) (see equation (8)) to extract frequency-domain features. Patches are sorted based on computed scores, with the highest- and lowest-frequency patches selected, resized, and processed through SRM (Fridrich and Kodovsky, 2012) to capture noise patterns. These frequency-based features are embedded using a ResNet-50 network, while high-level semantic features are extracted via ConvNeXt-based OpenCLIP. The outputs from both networks are concatenated and passed through a multi-layer perceptron (MLP) for final classification. This hybrid approach demonstrated better performance across benchmarks, including improvements on AIGCDetectBenchmark and GenImage (Zhu et al., 2024), while achieving competitive results on the challenging Chameleon dataset.
@@
Table 4. Evaluation of Detection Models from Frequency Domain Analysis Category on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Zhang et al. (Zhang et al., 2019) X X
Dzanic et al. (Dzanic et al., 2020) X X X
Liu et al. (Liu et al., 2022b)
Yousaf et al. (Yousaf et al., 2022) X
Bammey (Bammey, 2024) X
Meng et al. (Meng et al., 2024)
Li et al. (Li et al., 2024b)
Yan et al. (Yan et al., 2024)
2.4. Fingerprint Analysis Methods
@@
Fingerprint analysis in digital forensics refers to techniques that detect unique, traceable features left behind by devices or processes during image generation. Traditional approaches focused on detecting handcrafted features, including device fingerprints and postprocessing fingerprints. Device fingerprints, such as the photo-response nonuniformity (PRNU) pattern (Lukas et al., 2006), arise from manufacturing imperfections in imaging sensors, leaving a unique and stable mark on each captured image. Postprocessing fingerprints, on the other hand, originate from in-camera processing pipelines, including operations like demosaicking and compression, which embed specific patterns into images (Cozzolino and Verdoliva, 2019). While fingerprint analysis often overlaps with frequency domain methods, its primary focus is on extracting distinct fingerprints specific to generative models, such as GANs and diffusion models, rather than general spectral artifacts. The following methods are categorized under fingerprint analysis, as they aim to identify the unique traces of generative techniques.
2.4.1. GAN Fingerprint Analysis by Marra et al. (2019)
Marra et al. (Marra et al., 2019), building on the concept of PRNU (Lukas et al., 2006), introduced the idea that GANs, like digital cameras, imprint unique and stable fingerprints on their outputs, enabling source attribution and forensic analysis. The study focused on extracting GAN fingerprints using two models: CycleGAN (Zhu et al., 2017) and ProGAN (Karras, 2017). To estimate the fingerprint $F$, the residual signal $R_{i}$ between a GAN-generated image $X_{i}$ and its denoised version $f(X_{i})$ is computed:
$R_{i}=X_{i}-f(X_{i}),$
where $f(\cdot)$ is a denoising filter. The fingerprint is then estimated as the average residual over $N$ samples:
$\hat{F}=\frac{1}{N}\sum_{i=1}^{N}R_{i}.$
As $N$ increases, random noise diminishes, yielding a stable fingerprint.
To attribute an image to its GAN source, the correlation index between the image’s residual and the computed GAN fingerprint is evaluated. Experiments demonstrated that GAN fingerprints are distinct, with histogram analysis showing clear separations between fingerprints of different GANs. Additionally, training the same GAN architecture on different datasets produces unique fingerprints, highlighting the method’s generalizability for source attribution. This foundational work established GAN fingerprints as a robust tool for source attribution but noted the challenge of detecting cross-generator images without explicit training on all potential generator outputs.
2.4.2. GAN Fingerprint Attribution by Yu et al. (2019)
Yu et al. (Yu et al., 2019) proposed a framework for attributing GAN-generated images to their source by learning GAN fingerprints. This approach extends the work of Marra et al. (Marra et al., 2019) by systematically separating image fingerprints and model fingerprints instead of relying on hand-crafted features.
The image fingerprint $F_{I}$ is extracted using a reconstruction autoencoder, defined as the residual signal:
$F_{I}=I-R(I),$
where $R(I)$ is the reconstructed image generated by the autoencoder. The autoencoder is trained with a combination of pixel-wise $L_{1}$ and adversarial losses to ensure accurate reconstruction and artifact retention. The model fingerprint $F_{y}$ for each GAN $y\in Y$ is learned as a latent feature vector during training, encapsulating the architectural, dataset, and hyperparameter-specific traits unique to the GAN. Attribution is performed by computing the correlation between the image fingerprint $F_{I}$ and model fingerprints $F_{y}$ and maximizing this correlation over all potential sources. The process is supervised using a cross-entropy loss that encourages accurate attribution. This framework demonstrated improved accuracy in attributing images to their source GAN by effectively capturing both image-specific and model-specific traits, advancing the generalizability and robustness of GAN attribution methods.
2.4.3. Disentangling GAN Fingerprints by Yang et al. (2021)
@@
Yang et al. (Yang et al., 2021) proposed the GAN Fingerprint Disentangling Network (GFD-Net), advancing previous methods by focusing on disentangling content-irrelevant and GAN-specific fingerprints in synthetic images. Unlike earlier approaches that employ classification setups, GFD-Net integrates a generator, a discriminator, and an auxiliary classifier within an extended GAN framework to explicitly learn and separate these fingerprints. The generator adopts a U-Net architecture with an encoder-decoder structure. A classifier is added to the encoder’s output to predict the GAN source, enhancing the generator’s feature learning capability. The generator outputs a fingerprint $f$, which is added to a real image $x_{\text{real}}$ to produce a fingerprinted synthetic image $x_{\text{fp}}$:
@@
2.4.4. FingerprintNet: Synthesized Fingerprints for Generalized GAN Detection by Jeong: 2022
Jeong et al. (Jeong et al., 2022) proposed FingerprintNet, a framework aimed at generalizing GAN-generated image detection by synthesizing diverse GAN fingerprints. This approach addresses the challenge of detecting images from unseen GAN architectures without relying on GAN-specific training datasets. FingerprintNet employs an autoencoder-based fingerprint generator, incorporating random layer selection, multi-kernel deconvolution, and feature blending modules to create diverse and robust fingerprints. The generator is trained using a combination of reconstruction loss and similarity loss, ensuring that synthesized fingerprints accurately represent the characteristics of GAN-generated images.
For detection, FingerprintNet applies a Fast Fourier Transform (FFT) to the generated images to extract 2D spectra, where GAN fingerprints are more evident. A ResNet-50 (He et al., 2016) classifier is then used to distinguish real from generated images based on these fingerprint-highlighted spectra. To address dataset imbalance, the generator creates three fake images for each real image, and a mixed-batch strategy is applied during training to maintain balanced mini-batches. By synthesizing fingerprints and avoiding reliance on specific GAN datasets, FingerprintNet demonstrates improved generalization for detecting images from unseen GAN architectures, advancing robustness in GAN detection tasks.
2.4.5. Learning on Gradients by Tan: 2023
Tan et al. (Tan et al., 2023) proposed a novel detection framework, Learning on Gradients (LGrad), which leverages gradient-based representations as generalized artifacts. The framework begins by employing a transformative model, $T$, a pretrained CNN, to convert real and generated images into feature vectors. The gradient of $\text{sum}(l)$ with respect to the input image is then computed to capture generalized artifacts:
@@ 2.4.5. Learning on Gradients by Tan: 2023
Following this, a classification model, ResNet-50 (He et al., 2016), pretrained on the ImageNet dataset (Russakovsky et al., 2015), is trained on the computed gradients to learn underlying artifacts. Notably, $T$ remains fixed during gradient computation and is reused during inference to first extract gradients, which are then classified by the trained model. The authors highlight that training the classifier on computed gradients enables generalized learning of artifacts common across GAN models. The paper also compares the performance of various transformative models, including pretrained classifiers and GAN discriminators, showcasing their impact on detection performance. Readers are encouraged to refer to the main paper for a detailed analysis.
2.4.6. Data Augmentation in Fingerprint Domain By Wang: 2023 (Scaling & Mixup)
Wang et al. (Wang et al., 2023b) proposed a framework to enhance the generalizability of GAN-generated image detectors by augmenting synthetic data in the fingerprint domain. The method involves two key contributions: (1) extraction of fingerprints from synthetic images across different scenes using an encoder trained with Mean Squared Error (MSE) and adversarial losses, and (2) improved cross-GAN generalization through fingerprint perturbation. The framework utilizes an autoencoder trained on real images to extract residual fingerprints from GAN-generated images. These fingerprints, which represent GAN-specific artifacts, are made generalizable across scenes by incorporating a Gradient Reversal Layer (GRL) alongside MSE loss.

### v2

@@ 2.4.3 Two-Stream Convolutional Network for Fake Content Detection by Yousaf: 2022
Yousaf et al. Yousaf et al. (2022) proposed TwoStreamNet, a two-stream convolutional neural network designed to enhance the generalizability of fake visual content detection by jointly analyzing spatial and frequency features. The network comprises two main modules: the Spatial Stream and the Frequency Stream, which independently process input images and fuse their outputs at the classification stage to improve detection accuracy.
@@
The Frequency Stream captures frequency domain artifacts by first converting images to the YCbCr color space to decorrelate color channels. Discrete Fourier Transform (DFT) and Discrete Wavelet Transform (DWT) (see equation (3)) are applied to each channel. DFT decomposes image signals into real and imaginary components, while DWT captures both low- and high-frequency sub-bands. The resulting frequency features, represented as $H\times W\times 18$ feature maps, are processed using ResNet-50 He et al. (2016) to extract discriminative frequency patterns. The Spatial Stream focuses on spatial domain features by processing original RGB images augmented with Gaussian blur and JPEG compression, similar to the approach in Wang et al. (2020b). These augmented images are passed through a ResNet-50 network to extract spatial features. Finally, the outputs of the two streams are fused via probability averaging, ensuring equal contributions from spatial and frequency domains to the final decision. This combined framework highlights the importance of integrating frequency features for robust detection of synthesized visual content.
2.4.4 Synthbuster by Bammey: 2023
@@
Bammey Bammey (2024) introduced Synthbuster, a forensic method for detecting synthetic images generated by state-of-the-art diffusion models. The method begins by applying a cross-difference filter, originally defined by Chen et al. Chen and Hsu (2008), to highlight periodic frequency artifacts in synthetic images. The filter acts as a high-pass filter, generating residual images using the operation:
C(x,y)=|I(x,y)+I(x+1,y+1)I(x+1,y)I(x,y+1)|,C(x,y)=|I(x,y)+I(x+1,y+1)-I(x+1,y)-I(x,y+1)|,
@@
The Fast Fourier Transform (FFT) is then applied to the residual image, extracting spectral components corresponding to periodicities $0,2,4,$ and $8$ in both vertical and horizontal directions. The resulting features form a 135-dimensional magnitude peak vector for the RGB channels. These features are classified using a histogram-based gradient boosting tree classifier (HBGB) Ke et al. (2017), trained specifically to distinguish synthetic images from real ones.
@@
2.4.5 Meng’s Artifact Feature Purification: 2024
@@
Meng et al. Meng et al. (2024) addressed two key limitations in existing detection methods: poor generalization across generative models and limited effectiveness on images from diverse large-scale datasets. To overcome these challenges, they proposed the Artifact Purification Network (APN), which extracts generalizable artifact features through explicit and implicit purification processes. Explicit purification isolates artifact features in spatial and frequency domains by employing feature decomposition and frequency-band proposals to detect suspicious patterns. Implicit purification, guided by a classifier, further refines these features using mutual information estimation, enhancing the robustness of detection across various generators and datasets. This dual purification approach significantly improves generalization and detection accuracy.
2.4.6 An Image Transformation Perspective by Li: 2024
@@
Li et al. Li et al. (2024b) identified biases in existing detection methods, including weakened and overfitted artifact features, which limit their generalization capability. To address these challenges, the authors proposed the SAFE (Simple Preserved and Augmented Features) framework, designed to improve generalizable artifact detection by integrating effective transformations with preprocessing and augmentations. The framework employs several key strategies: (1) replacing conventional down-sampling with a crop operator (RandomCrop for training and CenterCrop for inference) to preserve local correlations and prevent artifact distortion; (2) applying invariant augmentations, such as ColorJitter and RandomRotation, to mitigate color discrepancies and irrelevant rotation-related features; (3) using a patch-based random masking strategy to enhance sensitivity to local image regions and subtle artifacts; and (4) incorporating Discrete Wavelet Transform (DWT) to extract high-frequency features critical for distinguishing synthetic from real images. The processed data was used to train a ResNet model He et al. (2016) with 1.44M parameters, achieving strong detection accuracy during testing. This approach demonstrated the effectiveness of integrating diverse transformations to improve detection robustness and generalization.
2.4.7 Yan’s AIDE Method for Synthetic Image Detection (2024)
@@
Yan et al. Yan et al. (2024) introduced the Chameleon dataset, a collection of AI-generated images designed to closely resemble real-world scenes, highlighting the limitations of existing detectors, which frequently misclassify these images as real. To address these challenges, the authors proposed AIDE (AI-generated Image Detector with Hybrid Features), a method that combines low-level frequency-based features with high-level semantic embeddings from CLIP. AIDE processes input images by dividing them into patches and applying discrete cosine transform (DCT) (see equation (4)) to extract frequency-domain features. Patches are sorted based on computed scores, with the highest- and lowest-frequency patches selected, resized, and processed through SRM Fridrich and Kodovsky (2012) to capture noise patterns. These frequency-based features are embedded using a ResNet-50 network, while high-level semantic features are extracted via ConvNeXt-based OpenCLIP. The outputs from both networks are concatenated and passed through a multi-layer perceptron (MLP) for final classification. This hybrid approach demonstrated better performance across benchmarks, including improvements on AIGCDetectBenchmark and GenImage Zhu et al. (2024), while achieving competitive results on the challenging Chameleon dataset.
Comparative Analysis of Frequency Domain Analysis Methods
Frequency-domain detection methods analyze the spectral signatures and periodic inconsistencies introduced by generative models, offering an orthogonal perspective to spatial feature analysis. Early work by Dzanic et al. Dzanic et al. (2020) demonstrated that GANs fail to reproduce high-frequency spectral content accurately, revealing that the Fourier spectrum decay parameters could effectively distinguish real from generated images. However, their approach was limited to GAN-generated data and relied on handcrafted frequency modeling. Liu et al. Liu et al. (2022b) shifted focus toward the intrinsic noise characteristics of real images, improving robustness by learning noise priors rather than relying solely on synthetic artifacts, thereby addressing over-dependence on generator-specific patterns. Yousaf et al. Yousaf et al. (2022) advanced this direction with TwoStreamNet, which fused spatial and frequency representations through parallel CNN streams, overcoming the modality isolation problem seen in prior methods that treated spatial and frequency cues independently. Bammey’s SynthBuster Bammey (2024) further extended frequency-based detection to diffusion models, addressing the prior gap in diffusion image forensics. By employing a cross-difference filter and spectral peak analysis, SynthBuster generalized frequency-based techniques beyond GANs and achieved robustness under compression. Building on these insights, Meng et al. Meng et al. (2024) introduced the Artifact Purification Network (APN), which directly tackled two recurring challenges—poor cross-model generalization and dataset dependency—by purifying spatial and frequency artifacts through explicit and implicit feature separation. Li et al. Li et al. (2024b) improved upon this with the SAFE framework, emphasizing preserved and augmented transformations, including DWT-based feature extraction, to strengthen generalization and mitigate overfitting observed in prior frequency-based CNNs. Yan et al. Yan et al. (2024) then unified these developments by combining frequency and semantic embeddings in AIDE, bridging low-level artifact cues with high-level CLIP features, and achieving superior robustness on challenging datasets like Chameleon, where earlier models failed due to limited semantic understanding. Collectively, these methods demonstrate a clear evolution—from handcrafted spectral analysis (Dzanic) and real-noise modeling (Liu) to dual-domain fusion (Yousaf), diffusion-oriented adaptation (Bammey), feature purification and augmentation (Meng, Li), and finally hybrid semantic–frequency reasoning (Yan). This trajectory reflects a shift from model-specific artifact detection toward unified, semantically informed, and cross-domain generalization in frequency-domain image forensics.
@@
Table 6: Evaluation of Detection Models from Frequency Domain Analysis Category on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Dzanic et al. Dzanic et al. (2020) X X X
Liu et al. Liu et al. (2022b)
Yousaf et al. Yousaf et al. (2022) X
Bammey Bammey (2024) X
Meng et al. Meng et al. (2024)
Li et al. Li et al. (2024b)
Yan et al. Yan et al. (2024)
2.5 Fingerprint Analysis Methods
@@
Fingerprint analysis in digital forensics refers to techniques that detect unique, traceable features left behind by devices or processes during image generation. Traditional approaches focused on detecting handcrafted features, including device fingerprints and postprocessing fingerprints. Device fingerprints, such as the photo-response nonuniformity (PRNU) pattern Lukas et al. (2006), arise from manufacturing imperfections in imaging sensors, leaving a unique and stable mark on each captured image. Postprocessing fingerprints, on the other hand, originate from in-camera processing pipelines, including operations like demosaicking and compression, which embed specific patterns into images Cozzolino and Verdoliva (2019). While fingerprint analysis often overlaps with frequency domain methods, its primary focus is on extracting distinct fingerprints specific to generative models, such as GANs and diffusion models, rather than general spectral artifacts. The following methods are categorized under fingerprint analysis, as they aim to identify the unique traces of generative techniques.
2.5.1 Disentangling GAN Fingerprints by Yang et al. (2021)
@@
Yang et al. Yang et al. (2021) proposed the GAN Fingerprint Disentangling Network (GFD-Net), advancing previous methods by focusing on disentangling content-irrelevant and GAN-specific fingerprints in synthetic images. Unlike earlier approaches that employ classification setups, GFD-Net integrates a generator, a discriminator, and an auxiliary classifier within an extended GAN framework to explicitly learn and separate these fingerprints. The generator adopts a U-Net architecture with an encoder-decoder structure. A classifier is added to the encoder’s output to predict the GAN source, enhancing the generator’s feature learning capability. The generator outputs a fingerprint $f$, which is added to a real image $x_{\text{real}}$ to produce a fingerprinted synthetic image $x_{\text{fp}}$:
@@
2.5.2 FingerprintNet: Synthesized Fingerprints for Generalized GAN Detection by Jeong: 2022
Jeong et al. Jeong et al. (2022) proposed FingerprintNet, a framework aimed at generalizing GAN-generated image detection by synthesizing diverse GAN fingerprints. This approach addresses the challenge of detecting images from unseen GAN architectures without relying on GAN-specific training datasets. FingerprintNet employs an autoencoder-based fingerprint generator, incorporating random layer selection, multi-kernel deconvolution, and feature blending modules to create diverse and robust fingerprints. The generator is trained using a combination of reconstruction loss and similarity loss, ensuring that synthesized fingerprints accurately represent the characteristics of GAN-generated images.
For detection, FingerprintNet applies a Fast Fourier Transform (FFT) to the generated images to extract 2D spectra, where GAN fingerprints are more evident. A ResNet-50 He et al. (2016) classifier is then used to distinguish real from generated images based on these fingerprint-highlighted spectra. To address dataset imbalance, the generator creates three fake images for each real image, and a mixed-batch strategy is applied during training to maintain balanced mini-batches. By synthesizing fingerprints and avoiding reliance on specific GAN datasets, FingerprintNet demonstrates improved generalization for detecting images from unseen GAN architectures, advancing robustness in GAN detection tasks.
2.5.3 Learning on Gradients by Tan: 2023
Tan et al. Tan et al. (2023) proposed a novel detection framework, Learning on Gradients (LGrad), which leverages gradient-based representations as generalized artifacts. The framework begins by employing a transformative model, $T$, a pretrained CNN, to convert real and generated images into feature vectors. The gradient of $\text{sum}(l)$ with respect to the input image is then computed to capture generalized artifacts:
@@ 2.5.3 Learning on Gradients by Tan: 2023
Following this, a classification model, ResNet-50 He et al. (2016), pretrained on the ImageNet dataset Russakovsky et al. (2015), is trained on the computed gradients to learn underlying artifacts. Notably, $T$ remains fixed during gradient computation and is reused during inference to first extract gradients, which are then classified by the trained model. The authors highlight that training the classifier on computed gradients enables generalized learning of artifacts common across GAN models. The paper also compares the performance of various transformative models, including pretrained classifiers and GAN discriminators, showcasing their impact on detection performance. Readers are encouraged to refer to the main paper for a detailed analysis.
2.5.4 Data Augmentation in Fingerprint Domain By Wang: 2023 (Scaling & Mixup)
Wang et al. Wang et al. (2023b) proposed a framework to enhance the generalizability of GAN-generated image detectors by augmenting synthetic data in the fingerprint domain. The method involves two key contributions: (1) extraction of fingerprints from synthetic images across different scenes using an encoder trained with Mean Squared Error (MSE) and adversarial losses, and (2) improved cross-GAN generalization through fingerprint perturbation. The framework utilizes an autoencoder trained on real images to extract residual fingerprints from GAN-generated images. These fingerprints, which represent GAN-specific artifacts, are made generalizable across scenes by incorporating a Gradient Reversal Layer (GRL) alongside MSE loss.